OpenStudy (anonymous):

I have to solve this in Algebra form: Pipes A, B, and C are connected to a swimming pool. Pipe A fills an empty pool in 45 minutes and pipe C empties a full pool in 35 minutes. Pipe B fills the pool. If all three pipes are left open then it takes them 30 minutes to fill an empty pool. How long would it take Pipe B alone to fill an empty pool?

OpenStudy (anonymous):

I don't think you wrote the whole question. One of the sentences just says "Pipe B fills the pool."

OpenStudy (anonymous):

Oh ignore that sentence! lol sorry about that, it's a typo.

OpenStudy (anonymous):

lets say that the pool is of size a liters. pipe a fills the pool at a rate of a/45 liters/min pipe b drains water at a rate of a/35 liters/min all 3 pipes fill the pool at a/30 liters/min the total effect of all pipes can be written as a/45 - a/35 + c = a/30 a/45 - a/35 - a/30 = -c 14a/630 - 18a/630 -21a/630 = -c c = 25a/630 or a/25.2 so pipe c fills the pool in 25.2 minutes lets just check that these numbers work with a pool that is 100 liters in size pipe a will flow at 100/45 pipe b will flow at 100/35 pipe c will flow at 100/25.2 the total flow will be 100/45+100/25.2-100/35 = 10/3 liters/minute so in 30 minutes it will be 30*(10/3) = 100 liters